OpenStudy (anonymous):

Please help me out. :( A body is moving at 40 m/s and its speed is decreasing uniformly at the rate of 5 m/s^2. Determine: a) its speed at the end of 6 sec? b) its average speed during 6 sec? c) the distance passed over in 6 sec?

OpenStudy (ahsome):

A) First, times the acceleration by the time \[5m/s^2*6=30\] Take away the speed by this number \[New Speed = OriginalSpeed - AboveNumber\] \[40-30=10\] The answer to A is 10

OpenStudy (anonymous):

i dont get this.. please explain.. hehe. :)

OpenStudy (ahsome):

Ok, first you know that his speed is decreasing at a rate of \[5ms^{-2}\] And you know this goes on for 6 seconds, as said in the question. Therefore, if you times the rate by 6, you can get how much the speed has decreased, here \[30ms^{-1}\] If you take away the total, here 40, from 30, you will get the current speed at 6 seconds, @itsmeeQUENNIE

OpenStudy (ahsome):

Does that help? @itsmeeQUENNIE

OpenStudy (anonymous):

ahhh.. yeah.. i got it. thank you for explaining it clearly. :)

OpenStudy (ahsome):

No problem, now to the next questions :)

OpenStudy (ahsome):

If, at any point, you are confused, just ask :)

OpenStudy (anonymous):

i used the formula.. Average Velocity = (v + u) / 2

OpenStudy (ahsome):

The equation for acceleration: \[Acceleration=\frac{v^{2}+v^{1}}2\] \[v^2=EndSpeed\] \[v^1=BeginningSpeed\] Sub in the values \[Acceleration=\frac{50+10}2\] \[Acceleration=30\]

OpenStudy (ahsome):

I meant average speed*

OpenStudy (ahsome):

3) \[AverageDistance=AverageSpeed*Time\] \[Distance=30*6\] \[Distance=180m\]

OpenStudy (ahsome):

That's it :)

OpenStudy (ahsome):

Anything you don't understand, @itsmeeQUENNIE?

OpenStudy (anonymous):

what is the answer in letter b?

OpenStudy (ahsome):

Answer for B) is 30m/s

OpenStudy (ahsome):

Using the rule you learn: \[AverageVelocity=\frac{v+u}{2}\]

OpenStudy (anonymous):

am, @Ahsome , why 50 ? why not 10 ? please explain. thankyou.

OpenStudy (ahsome):

@itsmeeQUENNIE What do you mean 50 and not 10? I used both in that equation

OpenStudy (anonymous):

ahm.. i mean.. why 50 is the final velocity?

OpenStudy (ahsome):

Sorry, I mean 10+50

OpenStudy (ahsome):

I put them in the wrong place ;)

OpenStudy (ahsome):

So it would be: \[AverageSpeed=\frac{10+50}{2}\]

OpenStudy (anonymous):

why 50 is the initial? how did you solve that?

OpenStudy (ahsome):

What am I doing. I meant 40 for the initial. My keyboard is stuck :(

OpenStudy (ahsome):

So, everytime I use 50, just put 40.

OpenStudy (ahsome):

@itsmeeQUENNIE Now do you understand?

OpenStudy (anonymous):

so, what's the final answer for letter b? :)

OpenStudy (ahsome):

To find Average Speed: \[AverageSpeed=\frac{40+10}{2}\] \[AverageSpeed=\frac{50}{2}\] \[AverageSpeed=25ms^{-2}\] To find Average Distance: \[AverageDistance=AverageSpeed∗Time\] \[Distance=25∗6\] \[Distance=150m\]

OpenStudy (anonymous):

in letter c, why did you use the answer in letter b? please explain it again. thankyou. :)

OpenStudy (anonymous):

u=40m/s a=-5m/s^2 (retardation) (a) t=6sec v=u+at v=40-5*6=40-30=10m/s (b) average speed \[=\frac{ u+v }{ 2 }=\frac{ 40+10 }{ 2 }=25 m/s\] (c) \[s=ut+\frac{ 1 }{ 2 }a t^2=40*6+\frac{ 1 }{ 2 }*(-5)*6^2=240-90=150 m\] or distance=average speed *time=25*6=150m

OpenStudy (ahsome):

Sure :) They asked you the distance they made in that six seconds. Remember the equation: \(Distance=Speed*Time\). Since the average speed is the overall speed he had in those 6 seconds, we use that. We don't use 50, because he didn't travle for \(50ms^{-1}\) for 6 seconds. He started at that speed, and slowly went down to \(10ms^{-1}\) at a rate of \(5ms^{-2}\). Instead of trying to use all that info, we just used the average speed. Kinda like how the average height is around 5Feet, 7Inches. Not many people ARE that height, but for mathematical reasons we can use it to cover all.

OpenStudy (ahsome):

@surjithayer I was WONDERING what you were typing there ;)

OpenStudy (ahsome):

There are two ways, either using the \(s=ut+12at^{2}\), as shown by @surjithayer. Or you can use the \(Distance=AverageSpeed*Time\)

OpenStudy (ahsome):

Hopefully that helps you, @itsmeeQUENNIE

OpenStudy (anonymous):

I guess.. i'll take the solution of @surjithayer .. i understand it easily without question on my mind. :) sorry @Ahsome . :( really really sorry.

OpenStudy (anonymous):

we can also use \[v^2-u^2=2as,\] \[\left( 10 \right)^2-\left( 40 \right)^2=2*(-5) s\] 100-1600=-10s -1500=-10s \[s=\frac{ -1500 }{ -10 }=150m\]

OpenStudy (ahsome):

@itsmeeQUENNIE It's OK :) Whatever floats your boat :)

OpenStudy (anonymous):

@surjithayer .. ah.. but im okay with your 1st solution..

OpenStudy (anonymous):

@Ahsome .. thank you so much for understanding.. :)

OpenStudy (ahsome):

No Problem

OpenStudy (anonymous):

you are welcome. i have only given you the various solutions for knowledge .

OpenStudy (anonymous):

thankyou. :) ^_^